solve the following system of linear equations and if possible express your answer as an ordered pair

y=2x+1
x+y=4

Answer by funmath(2933)

solve the following system of linear equations and if possible express your answer as an ordered pair
:
This is the substitution method:
:
L1) y=2x+1 Substitute 2x+1 in for y in L2) and solve for x.
L2) x+y=4
:
x+(2x+1)=4
x+2x+1-1=4-1
3x=3
3x/3=3/3
x=1 This is the x-coordinate, substitute 1 in for x in L1) and solve for y:
:
y=2(1)+1
y=2+1
y=3
The answer is: (x,y)=(1,3)
Happy Calculating!!!